OpenStudy (anonymous):

Write the complex number in the form a + bi. 3(cos 270° + i sin 270°)

OpenStudy (mertsj):

What is the cosine of 270?

OpenStudy (anonymous):

= 0

OpenStudy (mertsj):

What is the sin of 270?

OpenStudy (anonymous):

-1

OpenStudy (mertsj):

Substitute those values into 3(cos 270° + i sin 270°) and simplify

OpenStudy (anonymous):

i got 3?

OpenStudy (anonymous):

3i

OpenStudy (anonymous):

*

OpenStudy (mertsj):

0+3i in a + bi form

OpenStudy (anonymous):

these are my answer choices.. -3i 3i 3 -3

OpenStudy (anonymous):

so would it just be the 3i?

OpenStudy (mertsj):

3(0+-1i)=0-3i=-3i

OpenStudy (anonymous):

oh yeah duh lol thank you!! can you just tell me if my answer is correct on this other problem?

OpenStudy (mertsj):

What other problem?

OpenStudy (anonymous):

Find the product of z1 and z2 where z1 = 7(cos 40° + i sin 40°), and z2 = 6(cos 145° + i sin 145°). and for my answer i got .. 42(cos 185° + i sin 185°)

OpenStudy (anonymous):

?

OpenStudy (mertsj):

Yes. That is correct.

OpenStudy (anonymous):

ok thankyou!!

OpenStudy (mertsj):

yw